Calculate how much money $500,000 would provide in monthly income if the money were to last 35 years and the funds that were not withdrawn grew at a rate of 6% per year
Answer #1
Monthly income = Present value of monthly cash flow / Present value of annuity of 1
= $ 5,00,000.00 / 175.3802
= $ 2,850.95
working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005)^-420)/0.005 i = 6%/12 = 0.005
= 175.3802262 n = 35*12 = 420